WACC = 2%
Period Project 1 Project 2
0 -$150 -$60
1 $100 $50
2 $75 $35
3 $80
4
5
6

Based on the information in the table, what is the replacement chain NPV for Project 2?
Answer #1

As project 1 is of 3 years and 1 of 2 years the replacement will go on till project B reaches year 6. NPV is calculated below:

Project 2:

Year CF Discount Factor Discounted CF
0 $ -60.00 1/(1+0.02)^0= 1 1*-60= $   -60.00
1 $    50.00 1/(1+0.02)^1= 0.980392157 0.980392156862745*50= $     49.02
2 35-60 = $ -25.00 1/(1+0.02)^2= 0.961168781 0.961168781237985*-25= $   -24.03
3 $    50.00 1/(1+0.02)^3= 0.942322335 0.942322334547045*50= $     47.12
4 35-60 = $ -25.00 1/(1+0.02)^4= 0.923845426 0.923845426026514*-25= $   -23.10
5 $    50.00 1/(1+0.02)^5= 0.90573081 0.905730809829916*50= $     45.29
6 $    35.00 1/(1+0.02)^6= 0.887971382 0.887971382186192*35= $     31.08
NPV = Sum of all Discounted CF $ 172.50

In years 2 and 4 the project will be replaced

Replacement NPV for project 1 is given below:

Year CF Discount Factor Discounted CF
0 $ -150.00 1/(1+0.02)^0= 1 1*-150= $ -150.00
1 $    100.00 1/(1+0.02)^1= 0.980392157 0.980392156862745*100= $      98.04
2 $      75.00 1/(1+0.02)^2= 0.961168781 0.961168781237985*75= $      72.09
3 80-100 = $    -70.00 1/(1+0.02)^3= 0.942322335 0.942322334547045*-70= $    -65.96
4 $    100.00 1/(1+0.02)^4= 0.923845426 0.923845426026514*100= $      92.38
5 $      75.00 1/(1+0.02)^5= 0.90573081 0.905730809829916*75= $      67.93
6 $      80.00 1/(1+0.02)^6= 0.887971382 0.887971382186192*80= $      71.04
NPV = Sum of all Discounted CF $    401.48